# Hard View Selection for Self-Supervised Learning

### v4

@@ 5.2 Q2: Can a Manual Augmentation Policy be Inferred?
We further validated the previous result and assess an upper limit on the performance for predicting the hardest pair. By fitting a gradient boosting classification tree (Friedman, 2001), we predict the selected view pair conditioned on all SimSiam hyperparameter log data from Q1 except for the flag that indicates whether a view was selected. As training and test data, we used the logs from two seeds (each 300k samples) and the logs from a third seed, respectively. We also tuned hyperparameters on train/valid splits and applied a 5-fold CV. However, the resulting average test performance in all scenarios never exceeded 40%, indicating that it is indeed challenging to predict the hardness of views based on parameter-level data. This further supports our hypothesis that deriving a policy for controlling and increasing hardness based on geometric and appearance parameters is non-trivial and that such a policy must function on a per-sample basis and have access to the current model state as in HVS.
5.3 Q3: What are the Effects of Empowering the Adversary?
It is well known that adversarial learning can suffer from algorithmic instability (Xing et al., 2021), e.g. by giving an adversary too much capacity. We now further explore the space of adversarial capacity for selecting hard views to gain a better understanding of its impact on contrastive learning.
Optimizing an Adversarial Learner for the View Generation
In this experiment, we briefly explore adversarially learning a network to output the transformation matrix for view generation. We optimize a Spatial Transformer Network (STN) (Jaderberg et al., 2015) jointly with the DINO objective and a ViT-tiny. Using the inverse of the DINO loss, we train the STN to output the geometric transformation matrices (each 6 parameters) for generating views . We employ and test various transformation types (e.g., rotation, translation, etc.) that influence the capacity of the adversary. We discuss this experiment in greater detail in Appendix E and briefly summarize our findings here: applying an adversary in this way is very challenging due to the sensitivity of CL on the data augmentation and the dynamics of adversarial learning, often leading to either minor improvements or model collapses as we have observed when employing a too strong manual IoU policy in Q1.
Increasing the Number of Views
In our initial experiments, we explored variations in the number of sampled views $N$ with SimSiam and HVS. As can be seen in Fig. 7 in the appendix, while $N=8$ views still outperform the baseline in terms of linear evaluation accuracy, it is slightly worse than using $N=4$ views (-0.05% for 100 epochs and -0.14% for 200 epochs pretraining on linear evaluation). We interpret this result as the existence of a “sweet spot” in setting the number of views, where, in the limit, a higher number of views corresponds to approximating the adversarial learner described above. This would allow HVS to select progressively harder views that quickly results in very hard tasks and consequently in model collapse and performance deterioration.  [^1] For completeness, we also experimented with a cooperative HVS (see Appendix F.2 for details).
6 Conclusion
In this study, we presented HVS, a new data augmentation and learning strategy for Self-Supervised Learning designed to challenge pretrained models with more demanding samples. This method, while straightforward in its design, has proven to be a powerful tool, pushing the boundaries of the traditional random view generation in SSL. When combined with methods like DINO, SimSiam, iBOT, and SimCLR, HVS consistently showcased improvements in linear evaluation and a diverse set of transfer tasks. As the landscape of self-supervised learning continues to evolve, methods like HVS underscore the potential for further advancements in the field. By continuously challenging models with harder samples and refining augmentation techniques, we anticipate even greater strides in model performance and generalization in future research endeavors. While HVS introduces computational overhead due to its additional forward passes, our findings demonstrate that this overhead is significantly mitigated when using small resolution images in HVS without impeding its performance. This suggests that HVS is likely to continue to outperform baselines regardless of the training length.
Appendix A Examples Sampled by HVS
Figure 5: We depict row-wise ten example images from the ImageNet train set along with their sampled views with a finished, 100-epoch trained SimSiam ResNet50 model. Left: original image with the overlaid randomly sampled crops (colored dashed rectangles). Right: All views after applying resizing and appearance augmentations. The pair that is selected adversarially by HVS is highlighted in solid lines, eg. View 1 and View 4 in the first row.
Appendix B Training Loss
Figure 6: The training loss over 100 epochs. Comparing the DINO vanilla method with DINO + HVS. The spike and drop in the loss curve of DINO is caused by freezing the last layer in the first epoch which was proposed by the authors as a strategy to enhance downstream performance. For HVS we can only see a drop and no spike. We believe this is because HVS exposes the model to hard views from the beginning of training (i.e. the loss is immediately maximized).
Appendix C Effect of More Views on Linear Evaluation Performance
Appendix D Assessing the Importance of Metrics with fANOVA
To assess the importance of various metrics on the training loss, we apply fANOVA Hutter et al. (2014) on data that we logged during training with HVS. We used 300k samples that contain the following sampled parameters from the geometric and appearance data augmentation operations for each view: all random resized crop parameters (height and width of the original image, coordinates of crop corners and height and width of the crop), all Colorjitter (color distortion) strengths (brightness, contrast, saturation, hue), grayscale on/off, Gaussian blurring on/off, horizontal flip on/off, loss, and if the crop was selected or not. The metrics we chose are Intersection over Union (IoU), Relative Distance (sample-wise normalized distance of the center points of crop pairs), color distortion distance (the Euclidean distance between all four color distortion operation parameters, i.e. brightness, contrast, saturation, hue), and the individual color distortion parameters of the Colorjitter operation. As can be seen in Fig. 8, the metric with the highest predictive capacity on the loss is the IoU with an importance of 15.2% followed by brightness with 5.1%. The relative distance has an importance of 3.3%, the colorjitter distance 2.3%, the contrast 1.6%, the saturation 1.4%, hue 0.6%, and all parameters jointly 1.7%.
Figure 8: Application of fANOVA Hutter et al. (2014) on logged training data to determine metrics with high predictive capacity on the train loss.
Appendix E Adversarial Learner for the View Generation
@@
We use Spatial Transformer Network (STN) (Jaderberg et al., 2015) to allow generating views by producing 6D transformation matrices (allowing translating, rotating, shearing, scaling, affine transformations and combinations thereof) in a differentiable way since most common augmentations are not off-the-shelf differentiable. As described, we train it alongside the actual pretrained network using the same (inverted) objective. For our experiments we use DINO with multi-crop, i.e. 2 global and 8 local heads. As STN we use a small CNN with a linear layer for outputting the 10*6D transformation matrices. In this scenario, we use a ViT-tiny/16 with a 300 epoch pretraining on CIFAR10 with a batch size of 256. All other hyperparameters are identical to the ones reported in the DINO paper.
@@
As mentioned previously, the STN, without using auxiliary losses, starts zooming in and generating single-color views. To counteract this behavior, we experimented with different penalties on the transformation matrices produced by the STN. For instance, in order to limit the zooming pattern, we can use the determinants of the sub-matrices of the transformation matrix to penalize based on the area calculated and apply a regression loss (e.g. MSE). We refer to this type of penalty as Theta Crops Penalty (TCP). Additionally, we also restrict its parameters to stay within a sphere with different parameters for local and global crops. Next to determinant-based penalty losses, we also experimented with other penalty functions such as the weighted MSE between the identity and the current transformation matrix or penalties based on histograms of the input image and generated views after applying the transformation. To avoid strong uni-dimensional scaling behavior, we also implemented restricting scaling in a symmetric way (i.e. applied to both x and y dimensions) and refer to this as scale-sym.. We report our best results in Table 6 which are all TCP-based. As can be seen, no setting is able to outperform the baseline. Our best score was achieved with translation-scale-symmetric which is very similar to random cropping. When removing the symmetries in scaling, the performance drops further. Removing a constraint adds one transformation parameter and therefore one dimension. This can be seen as giving more capacity to the adversarial learner which in turn can make the task significantly harder. Similarly, when adding rotation, the performance drops further and in part drastically. This is on the one hand due to the penalties not being fully able to restrict the output of the STN. On the other, the task of extracting useful information from two differently rotated crops is even harder, and learning spatial invariance becomes to too challenging. All in all, we experienced two modes: either the STN is too restricted, leading to static output (i.e. independent of image content, the STN would produce constant transformation matrices) or the STN has too much freedom, resulting in extremely difficult tasks. See Fig. 10 for an example on the former behavior.
@@
Appendix F Additional Results
F.1 Object Detection and Instance Segmentation
For our object detection and instance segmentation analysis experiment, we report additional results in Table F.1 based on the AP metric. We use the C4 backbone variant (Wu et al., 2019) and finetuning with the 1x schedule.
Method Arch. VOC07+12 Object Det. COCO Instance Segm.
$\mathbf{AP_{all}}$ $\mathbf{AP_{50}}$ $\mathbf{AP_{75}}$ $\mathbf{AP_{mask}}$ $\mathbf{AP_{mask50}}$ $\mathbf{AP_{mask75}}$
SimSiam RN50 52.05 78.46 56.88 27.16 45.16 28.48
+ HVS RN50 52.91 79.06 57.55 27.93 46.51 29.36
Improvement +0.86 +0.60 +0.67 +0.77 +1.35 +0.88
DINO RN50 53.66 80.86 58.60 30.25 50.35 31.78
+ HVS RN50 52.86 80.51 58.21 30.02 50.39 31.52
Improvement +0.80 -0.36 -0.39 -0.23 +0.04 -0.27
Table 7:
F.2 Easy View Selection
@@
Appendix G Hyperparameters
G.1 Evaluations on ImageNet
G.1.1 DINO
For DINO, we report the ViT pretraining hyperparameters in Table 8. For ResNet-50, we use the same hyperparameters. Note, for HVS we limit the total number of comparisons to 128 across all heads. Linear evaluation is executed for 100 epochs and we use a learning rate of $0.001$, SGD optimizer (AdamW (Loshchilov & Hutter, 2019) during pretraining), a batch size of 1024, a momentum of $0.9$, and no weight decay.
@@ G.1.1 DINO
architecture vit_small epochs: 100
img_size 224 warmup_epochs: 10
patch_size 16 freeze_last_layer: 1
out_dim 65536 lr: 0.0005
norm_last_layer true min_lr: 1.0e-06
momentum_teacher 0.996 optimizer: AdamW
use_bn_in_head false weight_decay: 0.04
teacher_temp 0.04 weight_decay_end: 0.4
warmup_teacher_temp 0.04 global_crops_scale: 0.4, 1.0
warmup_teacher_temp_epochs 0 global_crops_size: 224
fp16 true local_crops_number: 8
batch_size 512 local_crops_scale 0.05, 0.4
clip_grad 3.0 local_crops_size: 96
drop_path_rate 0.1
Table 8: Pretraining ImageNet hyperparameters for the runs with DINO. For 300 epochs, we use a batch size of 1024.
G.1.2 SimSiam
In Table 9, we report the ResNet-50 pretraining hyperparameters. Linear evaluation is executed for 90 epochs (as reported by the SimSiam authors) and we use a learning rate of 0.1, LARS optimizer (You et al., 2017), a batch size of 4096, and no weight decay.
@@ G.1.2 SimSiam
batch_size 512
blur_prob 0.5
crops_scale 0.2, 1.0
crop_size 224
feature_dimension 2048
@@
predictor_dimension 512
weight_decay 0.0001
@@
G.1.3 SimCLR
We report the ResNet-50 pretraining hyperparameters for SimCLR in Table 10. Linear evaluation is executed for 90 epochs with a learning rate 0.1, SGD optimizer, batch size of 4096 and no weight decay.
@@ G.1.3 SimCLR
proj_hidden_dim 2048
out_dim 128
use_bn_in_head true
batch_size 4096
@@
warmup_epochs 10
zero_init_residual true
@@
G.2 Transfer to Other Datasets and Tasks
@@
G.3 Object Detection and Instance Segmentation
We have used the Detectron2 library (Wu et al., 2019) for object detection and instance segmentation. We followed the public codebase from MoCo (He et al., 2020) (like SimSiam) for all entries. Due to limited compute resources we changed the batch size. All parameters that differ from MoCo are reported in the Table 13. The pretrained models are finetuned end-to-end on the target datasets. All methods are based on 100-epoch pre-training on ImageNet.
VOC07+12 Object Det. Coce Inst. Segm.
Hyperparameter Value Value
batch_size 8 8
lr 0.01 0.01
steps 48000 90000
Table 13: Hyperparameters for VOC object detection and COCO instance segmentation.
Appendix H Computational Overhead of HVS
In Table 14, we compare the computational overhead factors across multiple baselines. The details of hardware and software used for this analysis are: NVIDIA RTX 3080, AMD R7 5800X, 32GB RAM, Ubuntu 22.04, PyTorch 2.0.1, CUDA 11.8. For DINO’s 2 global and 8 local views (default), applying HVS with nviews=2 sampled for each original view results in 4 global and 16 local views. Since considering all combinations would yield over 77k unique comparisons (4 over 2 times 16 over 8), to remain tractable, we limit the number of total comparisons to 128. Overall, reducing the image resolution can significantly reduce the time complexity of HVS. For instance, in the case of DINO ViT-S/16, the reduction leads to approximately 30% faster training times and as reported in the main paper, a 50% reduction allows retaining original performance.
Res. SimSiam (RN50) DINO (ViT)
100% x1.55 x2.15
75% x1.45 x1.68
50% x1.24 x1.56
25% x1.16 x1.5
Table 14: Time slowdown factors under different image resolutions for the selection in HVS.
While technically there can be a memory overhead with HVS, with the number of sampled views chosen in this paper, the backward pass of the methods that compute gradients only for the selected view pair still consumes more memory than the selection part of HVS (even for 8 sampled views in SimSiam). Note, that selection and the backward computation are never executed at the same time but sequentially.
We emphasize that the time overhead factors were measured without any optimization of HVS’ efficiency and, in our view, there are multiple ways to improve it. In our work, we opted for the easiest implementation possible to showcase that selecting harder views dependent on the model learning state can help boost performance in contrastive learning. Going forward with more compute-efficient HVS solutions, one could think of:

### v6

@@ F.1.2 Assessing the Difficulty of Predicting the Pair Selected by HVP
We further validate the previous result and assess an upper limit on the performance for predicting the hardest pair. By fitting a gradient boosting classification tree Friedman (2001), we predict the selected view pair conditioned on all SimSiam hyperparameter log data from Q1 except for the flag that indicates whether a view was selected. As training and test data, we used the logs from two seeds (each 300k samples) and the logs from a third seed, respectively. We also tuned hyperparameters on train/valid splits and applied a 5-fold CV. However, the resulting average test performance in all scenarios never exceeded 40%, indicating that it is indeed challenging to predict the hardness of views based on parameter-level data. This further supports our hypothesis that deriving a policy for controlling and increasing hardness based on geometric and appearance parameters is non-trivial and that such a policy must function on a per-sample basis and have access to the current model state as in HVP.
Appendix G Learning View Generation
G.1 Adversarial Learner
@@
In this experiment, we explore adversarially learning a network to output the transformation matrix for view generation. We use Spatial Transformer Network (STN) Jaderberg et al. (2015) to allow generating views by producing 6D transformation matrices (allowing translating, rotating, shearing, scaling, affine transformations, and combinations thereof) in a differentiable way since most common augmentations are not off-the-shelf differentiable. We optimize the STN jointly with the DINO objective and a ViT-tiny. We train it alongside the actual pretrained network using the same (inverted) objective. For our experiments, we use DINO with multi-crop, i.e. 2 global and 8 local heads. As STN we use a small CNN followed by a linear layer for outputting the 10*6D transformation matrices. In this scenario, we use a ViT-tiny/16 with a 300 epoch pretraining on CIFAR10 with a batch size of 256. All other hyperparameters are identical to the ones reported in the DINO paper.
@@
The STN, without using auxiliary losses, starts zooming in and generating single-color views. To counteract this behavior, we experimented with different penalties on the transformation matrices produced by the STN. For instance, in order to limit the zooming pattern, we can use the determinants of the sub-matrices of the transformation matrix to penalize based on the area calculated and apply a regression loss (e.g. MSE). We refer to this type of penalty as Theta Crops Penalty (TCP). Additionally, we also restrict its parameters to stay within a sphere with different parameters for local and global crops. Next to determinant-based penalty losses, we also experimented with other penalty functions such as the weighted MSE between the identity and the current transformation matrix or penalties based on histograms of the input image and generated views after applying the transformation. To avoid strong uni-dimensional scaling behavior, we also implemented restricting scaling in a symmetric way (i.e. applied to both x and y dimensions) and refer to this as scale-sym.. We report our best results in Table 5 which are all TCP-based. As can be seen, no setting is able to outperform the baseline. Our best score was achieved with translation-scale-symmetric which is very similar to random cropping. When removing the symmetries in scaling, the performance drops further. Removing a constraint adds one transformation parameter and therefore one dimension. This can be seen as giving more capacity to the adversarial learner which in turn can make the task significantly harder. Similarly, when adding rotation, the performance drops further and in part drastically. This is on the one hand due to the penalties not being fully able to restrict the output of the STN. On the other, the task of extracting useful information from two differently rotated crops is even harder, and learning spatial invariance becomes too challenging. All in all, we experienced two modes: either the STN is too restricted, leading to static output (i.e. independent of image content, the STN would produce constant transformation matrices) or the STN has too much freedom, resulting in extremely difficult tasks. See Fig. 10 for an example of the former behavior.
@@
G.2 Cooperative Learner
@@
Appendix H Additional Results
H.1 Object Detection and Instance Segmentation
Here we provide more detailed results on object detection and instance segmentation, shown in Tab. 6 and Tab. 7 respectively. We followed iBOT’s default configurations which employed Cascade Mask R-CNN as the task layer.
Method Arch. Object Detection Instance Segmentation
$\mathbf{AP^{b}}$ $\mathbf{AP^{b}_{50}}$ $\mathbf{AP^{b}_{75}}$ $\mathbf{AP^{m}}$ $\mathbf{AP^{m}_{50}}$ $\mathbf{AP^{m}_{75}}$
iBOT ViT-S/16 47.00 66.13 50.63 40.67 63.10 43.37
+ HVP ViT-S/16 47.27 66.50 50.90 40.90 63.50 43.83
Improvement +0.27 +0.37 +0.27 +0.23 +0.40 +0.47
DINO ViT-S/16 46.50 65.90 50.30 40.43 62.83 43.27
+ HVP ViT-S/16 47.07 66.37 50.63 40.80 63.37 43.87
Improvement +0.57 +0.47 +0.33 +0.37 +0.53 +0.60
Table 6: Additional object detection and instance segmentation results on the COCO dataset. The ViT-S/16 models were pretrained for 100 epochs.
Method Arch. Object Detection Instance Segmentation
$\mathbf{AP^{b}}$ $\mathbf{AP^{b}_{50}}$ $\mathbf{AP^{b}_{75}}$ $\mathbf{AP^{m}}$ $\mathbf{AP^{m}_{50}}$ $\mathbf{AP^{m}_{75}}$
iBOT ViT-S/16 47.60 66.80 51.33 41.10 63.63 44.07
+ HVP ViT-S/16 48.03 67.13 51.73 41.50 64.23 44.40
Improvement +0.43 +0.33 +0.40 +0.40 +0.60 +0.33
DINO ViT-S/16 47.27 66.60 51.00 41.00 63.63 44.03
+ HVP ViT-S/16 47.50 67.00 51.33 41.17 64.00 44.13
Improvement +0.23 +0.40 +0.33 +0.17 +0.37 +0.10
Table 7: Additional object detection and instance segmentation results on the COCO dataset. The ViT-S/16 models were pretrained for 300 epochs.
Appendix I Hyperparameters
I.1 Evaluations on ImageNet
I.1.1 DINO
For DINO, we report the ViT pretraining hyperparameters in Table 8 (ViT-S) and Table 9 (ViT-B) which are the original ones as reported by the authors. Note, for HVP we limit the total number of comparisons to 128 across all heads. Linear evaluation is executed for 100 epochs and we use a learning rate of $0.00075$, SGD optimizer (AdamW Loshchilov & Hutter (2019) during pretraining), a batch size of 1024, a momentum of $0.9$, and no weight decay (as reported by the authors).
@@ I.1.1 DINO
architecture vit-small epochs: 100
img-size 224 warmup-epochs: 10
patch-size 16 freeze-last-layer: 1
out-dim 65536 lr: 0.0005
norm-last-layer true min-lr: 1.0e-06
momentum-teacher 0.996 optimizer: AdamW
use-bn-in-head false weight-decay: 0.04
teacher-temp 0.04 weight-decay-end: 0.4
warmup-teacher-temp 0.04 global-crops-scale: 0.4, 1.0
warmup-teacher-temp-epochs 0 global-crops-size: 224
fp16 true local-crops-number: 8
batch-size 512 local-crops-scale 0.05, 0.4
clip-grad 3.0 local-crops-size: 96
drop-path-rate 0.1
Table 8: Pretraining ImageNet hyperparameters for the runs with DINO ViT-S/16. For 300 epochs, we use a batch size of 1024.
Hyperparameter Value Hyperparameter Value
architecture vit-base epochs: 400
img-size 224 warmup-epochs: 10
patch-size 16 freeze-last-layer: 3
out-dim 65536 lr: 0.00075
norm-last-layer true min-lr: 2.0e-06
momentum-teacher 0.996 optimizer: AdamW
use-bn-in-head false weight-decay: 0.04
teacher-temp 0.07 weight-decay-end: 0.4
warmup-teacher-temp 0.04 global-crops-scale: 0.25, 1.0
warmup-teacher-temp-epochs 50 global-crops-size: 224
fp16 false local-crops-number: 10
batch-size 1024 local-crops-scale: 0.05, 0.25
clip-grad 0.3 local-crops-size: 96
drop-path-rate 0.1
Table 9: Pretraining ImageNet hyperparameters for the runs with DINO ViT-B/16.
I.1.2 SimSiam
In Table 10, we report the ResNet-50 pretraining hyperparameters. Linear evaluation is executed for 90 epochs (as reported by the SimSiam authors) and we use a learning rate of 0.1, LARS optimizer You et al. (2017), a batch size of 4096, and no weight decay.
@@ I.1.2 SimSiam
batch-size 512
blur-prob 0.5
crops-scale 0.2, 1.0
crop-size 224
feature-dimension 2048
@@
predictor-dimension 512
weight-decay 0.0001
@@
I.1.3 SimCLR
We report the ResNet-50 pretraining hyperparameters for SimCLR in Table 11. Linear evaluation is executed for 90 epochs with a learning rate of 0.1, SGD optimizer, batch size of 4096, and no weight decay.
@@ I.1.3 SimCLR
proj-hidden-dim 2048
out-dim 128
use-bn-in-head true
batch-size 4096
@@
warmup-epochs 10
zero-init-residual true
@@
I.2 Transfer to Other Datasets and Tasks
@@
I.3 Object Detection and Instance Segmentation
Our experiments utilized Open MMLab’s detection library Chen et al. (2019) for object detection and instance segmentation on COCO Lin et al. (2014). We followed iBOT’s default configuration.
Obj. Det. & Inst. Segm. on COCO
Hyperparameter Value
epochs 12
batch-size 32
lr 0.02
Table 14: Hyperparameters object detection and instance segmentation on COCO.
Appendix J Computational Overhead of HVP
The additional forward passes that HVP introduces for the selection phase increase the time complexity of the individual baseline methods. Several possible approaches exist to mitigate this overhead, one of which is to alternate between the vanilla and HVP training step. We measured the overhead factors for different alternating frequencies (i.e., after how many training steps we use the hard views from HVP; we refer to this as step) for SimSiam, DINO, and iBOT which we report in Table 15. Below, we propose additional ways that may allow using hard views more efficiently.
Step
SimSiam
DINO
(ViT-S/16)
DINO
iBOT
(ViT-S/16)
1 (HVP) x1.64 x2.21 x2.01 x2.13
2 x1.38 x1.61 x1.56 x1.57
3 x1.32 x1.43 x1.42 x1.38
4 x1.29 x1.34 x1.35 x1.29
Table 15: Slowdown factors for HVP and the alternating training method, where step refers to the interval at which HVP is applied during training (i.e., step=1 refers to full HVP and step=3 indicates that HVP is approx. used 33% of the training time). Measurements are averages across 3 seeds.
As an alternative to the alternating training, we also experimented with a 50% image resolution reduction for the selection phase but observed that the final performance was negatively affected by it or that baseline performance could not be improved.
The details of hardware and software used for this analysis are: one single compute node with 8 NVIDIA RTX 2080 Ti, AMD EPYC 7502 (32-Core Processor), 512GB RAM, Ubuntu 22.04.3 LTS, PyTorch 2.0.1, CUDA 11.8. For DINO’s 2 global and 8 local views (default), applying HVP with nviews=2 sampled for each original view results in 4 global and 16 local views. Since considering all combinations would yield over 77k unique comparisons ($\binom{4}{2}\times\binom{16}{8}$), to remain tractable, we limit the number of total comparisons to 64. For training experiments that exceed the limit of 8x RTX 2080 Ti, we apply gradient accumulation.
While technically there can be a memory overhead with HVP, with the number of sampled views chosen in this paper, the backward pass of the methods that compute gradients only for the selected view pair still consumes more memory than the selection part of HVP (even for 8 sampled views in SimSiam). Note, that selection and the backward computation are never executed at the same time but sequentially.
We emphasize that further ways exist to optimize HVP’s efficiency which remain to be explored. For instance: